In Mordica Company, total materials costs are $33,000, and total conversion costs are $57,000. Equivalent units of production are materials 10,000 and conversion costs 12,000.

Compute the unit costs for materials and conversion costs. (Round answers to 2 decimal places, e.g. 2.25.)

Materials cost per unit

$enter a dollar amount rounded to 2 decimal places

Conversion cost per unit

$enter a dollar amount rounded to 2 decimal places

eTextbook and Media

  

  

Compute total manufacturing costs. (Round answer to 2 decimal places, e.g. 2.25.)

Total manufacturing cost per unit

$enter a dollar amount of total manufacturing cost per unit rounded to 2 decimal places

eTextbook and Media
Answer #1

Part-1

Material Cost per Unit = Totla Material cost/ Equivalent Unit of Production

$33000/10000= $3.30 per Unit
Conversion Cost per Unit = Totla Conversion cost/ Equivalent Unit of Production
=57000/12000= $4.75 per Unit
Part-2
Totl Manufacturing Cost= Material Cost per Unit+ Conversion Cost per Unit
=3.3 + 4.75= $8.05 per Unit